### v2

@@ 4.1 Tests based on simulated data
In this section, we compare the three different strategies: RDI, ADI, and ARDI, all utilizing the same core algorithm, IPCA. We consistently employed the same evaluation pipeline as presented in Juillard et al. (2023), which was previously used to compare three different algorithms for processing data sets using ADI alone. This test pipeline consists of a total of 60 test data sets composed of five different disk morphologies, injected at three different contrast levels ($10^{-3}$, $10^{-4}$, and $10^{-5}$), into four different observing ADI sequences of stars without any known circumstellar signal, reflecting different observing conditions. The test datasets are available on Zenodo. [^1] Datasets available at https://zenodo.org/records/11442267
@@
Figure 8: Four examples of IPCA disk estimations obtained using RDI, ADI, and ARDI strategies, for test dataset where ARDI performed significantly better than ADI and RDI. From top to bottom, the test dataset are cube #2 with disk A injected at a contrast of $10^{-3}$, cube #2 with disk E at a contrast of $10^{-4}$, cube #1 with disk D at a contrast of $10^{-4}$, and cube #4 with disk C at a contrast of $10^{-4}$. From left to right, the images display the injected disk and the results using RDI, ADI, and ARDI, utilizing the IPCA algorithm. These results were selected as they demonstrate ARDI’s capability to compensate for the individual limitations of each strategy compared to using each method individually, in a variety of scenarios in terms of observing conditions, disk morphologies, and contrasts.
@@ 4.2 Discussion of the test results
In our second series of tests, we used the IPCA algorithm with three different selections of reference stars with different levels of quality according to PCC (see Fig. 5). The data set #4 (very stable) seems more sensitive to the selection of reference frames, while data sets showing more speckle variability (#3, #2, #1) seem less sensitive. The mean PCC calculated for a non-optimal selection is significantly lower for cube #4 (mean PCC $\mu<0.6$) compared to other cubes (mean PCC $\mu\sim 0.8-0.5$). As a result, the estimates produced when using these sub-optimal references significantly under-performed, with SSIM hardly reaching 0.8 in most cases, while for other cubes, sub-optimal references could reach similar SSIM as the optimal and semi-optimal selections (see Fig. 7). Most interestingly, the three different selections achieved very similar results in terms of SSIM in many cases. In more than 40% of cases, non-optimal references achieved a higher SSIM than the optimal selection. The cases where the references contained 25% of optimal references are very close to an optimal selection, even for a contrast of $10^{-5}$. This suggests that selecting high-quality references, as assessed via PCC, has limited impact on the results in the final post-processed images using IPCA with ARDI. Instead, the outcomes appear to be more sensitive to the stability of the data set.
Figure 9: Gallery of protoplanetary disks data sets, originally described by Ren et al. (2023), re-processed here with IPCA-ARDI. The images have dimensions of $162\times 162$ pixels ($\sim 2\arcsec$) with logarithmic scale color bars. Images marked by a yellow star are those that were not presented in Ren et al. (2023) due to the low quality of the produced disk estimation, but that we still consider as a detection in our IPCA-ARDI results. Images marked with a red name are those that are not considered clear detection with IPCA-ARDI.
@@ 5 Demonstration on protoplanetary disk data sets
To test our algorithm on real observations, we used the preprocessed data sets on protoplanetary disks from Ren et al. (2023). The sample consists of 48 data sets including 29 unique different young stars hosting protoplanetary disks with known substructures previously observed in scattered light. Among them, four were reported to have exoplanet candidates (or claims thereof) from high-contrast imaging: HD 100546 (Quanz et al. 2013, 2015; Currie et al. 2014), HD 169142 (Biller et al. 2014; Reggiani et al. 2014; Gratton et al. 2019; Hammond et al. 2023), LkCa 15 (Kraus & Ireland 2012; Sallum et al. 2015), and MWC 758 (Reggiani et al. 2018; Wagner et al. 2019, 2023). Additionally, localised deviations from Keplerian rotation possibly associated with the presence of planets were reported in HD 163296 (Izquierdo et al. 2022; Pinte et al. 2020, 2018) and HD 97048 (Pinte et al. 2019). All observations were made in the broad-band $Ks$ filter using the IRDIS camera of VLT/SPHERE. These disks were previously presented by Ren et al. (2023) in both polarized light using the IRDAP (van Holstein et al. 2020) pipeline for polarimetric data processing, and in total intensity using RDI through the DI-sNMF algorithm. The references were acquired via the “star-hopping” mode (Wahhaj et al. 2021), which means that the telescope is regularly pointed at a nearby reference star during data acquisition, efficiently alternating science and reference observations with minimal overhead. This observing strategy ensures maximal similarity between the respective PSFs.
@@ 5.1 Inspecting protoplanet claims in individual systems
Reggiani et al. (2014) and Biller et al. (2014) independently observed a PLF in $L^{\prime}$ band using independent ADI data sets from VLT/NACO processed with PCA. Better images of the disk obtained with SPHERE suggested that this candidate at $0\aas@@fstack{\prime\prime}11$ separation may rather trace filtered signal from the inner ring of the disk (e.g. Ligi et al. 2019). Later, another candidate was proposed using observations in polarized light from 2015, along with two ASDI observations in 2017 and 2019 using the IRDIS and multiple IFS observations between 2015 and 2019 (Gratton et al. 2019; Hammond et al. 2023). However, detecting the candidate with ASDI required suppressing most of the disk signal. The observations were conducted in $J$ and $YJH$ bands and the position of the candidate between 2015 and 2019 appeared to follow a Keplerian motion. The signal was interpreted as a gap-clearing Jovian-mass protoplanet surrounded by a circumplanetary disk or envelope. Studies by Garg et al. (2021) and Law et al. (2023) using ALMA observations revealed several distinct chemical signatures that demonstrate a compelling link to ongoing giant planet formation. Furthermore, compact CO J=2–1 and CO J=3–2 emission counterparts were observed that coincide with the location of the candidate planet. Yet, no PLF appears in polarized light nor in total intensity for both the DI-sNMF and IPCA-ARDI of the $K$-band images presented here, at its predicted location at the epoch of the Ks observations considered in this work (2021) assuming Keplerian motion. In Appendix F.2, we attempt to infer the detectability of the candidate by injecting a fake planet at the same separation, using the estimated contrast in Hammond et al. (2023). In this test, we processed the data set with IPCA-ARDI as well as plain PCA-ADI. In addition, we provide the contrast curve obtained via PCA-ADI in Fig. 18. . Our $5\sigma$ detection limit of $1\times 10^{-4}$ at the expected location of the companion translates into $2M_{\text{Jup}}$ according to the ATMO2020 model, assuming no extinction. In comparison, Hammond et al. (2023) report a $YJH$ contrast of $\sim 1.5\times 10^{-5}$. Nevertheless, the detection of polarized intensity colocated with the protoplanet suggests the measured $YJH$ light is dominated by scattered light, making the mass conversion using atmospheric models most probably irrelevant. Assuming similar expected contrast in $Ks$ band as in $YJH$, our senitivity limits do not allow us to confirm or refute the candidate.
@@
Using CO line observations, Pinte et al. (2018) initially reported a local deviation from Keplerian velocity, suggesting the presence of a massive body at approximately 260 au from the star. Subsequently, Izquierdo et al. (2022) re-observed the first candidate proposed by Pinte et al. (2018), along with a second localized deviation from Keplerian rotation, potentially associated with the presence of a giant planet at 94 au. The first candidate proposed by Pinte et al. (2018) and Izquierdo et al. (2022) at 260 au does not fit within the field of view of the images presented in Figs. 11 and 9. At these separations, the motion of the candidates between the date of the first detection and the new dataset is negligible. We show another post-processed image for the same data set, with an uncropped field-of-view in Appendix F.4 as well as contrast curve with PCA-ADI. No PLF is observed at the locations of the two proposed sources in these direct imaging observations with both IPCA-ARDI and PCA-ADI, despite the estimated sensitivity reaching down to $3\times 10^{-4}$ in contrast (i.e, mass limit of $1M_{J}$) for the candidate at 260 au, and $6\times 10^{-6}$ in contrast (i.e., mass limit of $4M_{J}$) at 94 au (see details in Appendix F.4).
@@ 6 Conclusion
We also explored the problem of parameters optimization for IPCA algorithms. While a good choice of parameters is important, we found that multiple sets of parameters lead to similar results. Nevertheless, we identified different scenarios and matched them with specific parametrizations to provide optimal results in each case. For the brightest disks with a stable speckle field, it is better to start the iterative process at rank one and have multiple iterations per rank before increasing the rank (e.g., CPD-68 1894 or HD 36112 data presented in Sect. 5). This ensures that most of the self-subtraction effects have been mitigated before loosening the model of the speckle field by increasing the rank. Conversely, when the disk is faint relative to the speckle, it is necessary to start at a high rank and limit the number of iterations per rank to avoid bright artifacts stemming from non-captured speckle field components being propagated in the next iterations (e.g., V1094 Sco or GM Aur data sets). Regarding the convergence of the iterative process, we acknowledge the difficulty in building a reliable stopping criterion, which is a common problem in computer science (Nocedal et al. 2002). It is particularly challenging in our case of application, as determining where to stop the iterative process involves choosing a good compromise between noise subtraction (non-propagation) and filtering/recovery of extended signals (convergence), and also because different parts of the disk converge at different rates. Nevertheless, observing the convergence curves remains an interesting indicator of the optimal iteration to be reached.
@@
While a few planet candidates were reported via direct imaging among the sampled young stars, we did not re-observe any of these candidates. This is at least partly due to limitations in the achieved sensitivity in the considered filter ($Ks$), which renders the detection of these companions challenging or unfeasible. While confirmed detections of protoplanets within disks remain rare, with only a handful reported in the literature (e.g., Keppler et al. 2018; Haffert et al. 2019), studying young systems in their formative stages embedding protoplanetary disks housing nascent planets is crucial for understanding the intricate process of planetary system formation. The application of IPCA-ARDI holds promise in providing valuable insights by preserving both the structures of the disk and clear images of planets.
@@ Appendix A Local convergence of IPCA-ARDI
Figure 12: Evolution of the IPCA with ARDI estimate through the iterations. The data set used is the same as in Figure 3. The first grid of plots (top) displays the final estimation segmented into patches of size $15\times 15$ pixels. The second grid of plots (bottom) corresponds to the gradient computed within these patches according to Equation 4. The red lines indicate the $g_{\rm tol}$ threshold at ±$\pm 10^{-2}$. If the threshold is reached, the curves are displayed in green, and a black dot indicates at which iteration the signal has converged. Otherwise, the curves are displayed in blue.
In this appendix, we study the evolution of the integrated flux into different patches of the image through the IPCA iterative process when using ARDI. Figure 12 shows that the brightest part of the disk converges faster than the rest of the image. We can also observe that the parts of the image where no signal is present do not seem to converge. It is important to note that the observed non-convergence on the outer parts of the image where there is no disk signal relates to the expression of the gradient. Indeed, from Eq. 4 it is clear that for an estimate $d_{i}$ that tends towards 0, the quantity $\frac{d_{i}-d_{i+1}}{d_{i}}$ becomes more sensitive to minor variations.
@@ Appendix C Comparison of results for different metrics
Figure 14: Same as Fig. 6 for the four different metrics, respectively from top to bottom: Spearman’s rank, Pearson correlation, Euclidean distance, and SAD.
@@ Appendix D Results of disk estimations
Table 2: Table summarizing the IPCA parameters used to re-process each data set from Ren et al. (2023) in Sect. 5. For each data set, this table provides the starting rank and the number of iterations per rank. Regarding the number of iterations per rank, “incr” stands for incremental and indicates that the number of iterations increases by one at each rank (e.g., one iteration is performed at the starting rank $q$, two at rank $q+1$, and so on). Additionally, we also provide the number of iterations performed to obtain the final frame selected from visual assessment (“final iter.”), the rank that was reached at this iteration (“final rank”), and the mean and standard deviation of the PCC between science images and references.
@@ F.1 HD 36112 (MWC 758)
Figure 15: Results of IPCA (left) and 3-rank PCA with ADI-only (right) for the VLT/SPHERE data set of HD 36112 (Ren et al. 2023). Three simulated companions at a separation of $0\aas@@fstack{\prime\prime}6$ have been injected into the cube, with contrasts of $8\times 10^{-6}$, $1\times 10^{-5}$, and $2\times 10^{-5}$. White lines indicate their positions. Two squares are displayed to mark the locations of the two claimed protoplanets: a red square at $0\aas@@fstack{\prime\prime}1$, indicating the candidate proposed by Reggiani et al. (2018), and a yellow square at $0\aas@@fstack{\prime\prime}6$, referring to the candidate proposed by Wagner et al. (2019).
Figure 16: Contrast curve at $5\sigma$ confidence for the HD 36112 data set, processed using rank-3 PCA with ADI-only.

### v3

@@ 4.1 Tests based on simulated data
In this section we compare the three different strategies: RDI, ADI, and ARDI, all utilizing the same core algorithm, IPCA. We consistently employed the same evaluation pipeline as presented in Juillard et al. (2023), which was previously used to compare three different algorithms for processing datasets using ADI alone. This test pipeline consists of a total of 60 test datasets composed of five different disk morphologies, injected at three different contrast levels ($10^{-3}$, $10^{-4}$, and $10^{-5}$), into four different observing ADI sequences of stars without any known circumstellar signal, reflecting different observing conditions. The test datasets are available on Zenodo. [^1] Datasets available at https://zenodo.org/records/11442267
@@
Figure 8: Four examples of IPCA disk estimations obtained using RDI, ADI, and ARDI strategies, for a test dataset where ARDI performed significantly better than ADI and RDI. From top to bottom, the test dataset are cube #2 with disk A injected at a contrast of $10^{-3}$, cube #2 with disk E at a contrast of $10^{-4}$, cube #1 with disk D at a contrast of $10^{-4}$, and cube #4 with disk C at a contrast of $10^{-4}$. From left to right, the images display the injected disk and the results using RDI, ADI, and ARDI, utilizing the IPCA algorithm. These results were selected as they demonstrate ARDI’s capability to compensate for the individual limitations of each strategy compared to using each method individually, in a variety of scenarios in terms of observing conditions, disk morphologies, and contrasts.
@@ 4.2 Discussion of the test results
In our second series of tests, we used the IPCA algorithm with three different selections of reference stars with different levels of quality according to PCC (see Fig. 5). The dataset #4 (very stable) seems more sensitive to the selection of reference frames, while datasets showing more speckle variability (#3, #2, #1) seem less sensitive. The mean PCC calculated for a nonoptimal selection is significantly lower for cube #4 (mean PCC $\mu<0.6$) compared to other cubes (mean PCC $\mu\sim 0.8-0.5$). As a result, the estimates produced when using these suboptimal references significantly under-performed, with SSIM hardly reaching 0.8 in most cases, while for other cubes, suboptimal references could reach similar SSIM as the optimal and semi-optimal selections (see Fig. 7). Most interestingly, the three different selections achieved very similar results in terms of SSIM in many cases. In more than 40% of cases, nonoptimal references achieved a higher SSIM than the optimal selection. The cases where the references contained 25% of optimal references are very close to an optimal selection, even for a contrast of $10^{-5}$. This suggests that selecting high-quality references, as assessed via PCC, has limited impact on the results in the final post-processed images using IPCA with ARDI. Instead, the outcomes appear to be more sensitive to the stability of the dataset.
Figure 9: Gallery of protoplanetary disks datasets, originally described by Ren et al. (2023), and re-processed here with IPCA-ARDI. The images have dimensions of $162\times 162$ pixels ($\sim 2\arcsec$) with logarithmic scale color bars. Images marked by a yellow star are those that were not presented in Ren et al. (2023) due to the low quality of the produced disk estimation, but that we still consider as a detection in our IPCA-ARDI results. Images marked with a red name are those that are not considered clear detection with IPCA-ARDI.
@@ 5 Demonstration on protoplanetary disk datasets
To test our algorithm on real observations, we used the preprocessed datasets on protoplanetary disks from Ren et al. (2023). The sample consists of 48 datasets including 29 unique different young stars hosting protoplanetary disks with known substructures previously observed in scattered light. Among them, four were reported to have exoplanet candidates (or claims thereof) from high-contrast imaging: HD 100546 (Quanz et al. 2013, 2015; Currie et al. 2014), HD 169142 (Biller et al. 2014; Reggiani et al. 2014; Gratton et al. 2019; Hammond et al. 2023), LkCa 15 (Kraus & Ireland 2012; Sallum et al. 2015), and MWC 758 (Reggiani et al. 2018; Wagner et al. 2019, 2023). Additionally, localized deviations from Keplerian rotation possibly associated with the presence of planets were reported in HD 163296 (Izquierdo et al. 2022; Pinte et al. 2020, 2018) and HD 97048 (Pinte et al. 2019). All observations were made in the broad-band $Ks$ filter using the IRDIS camera of VLT/SPHERE. These disks were previously presented by Ren et al. (2023) in both polarized light using the IRDAP (van Holstein et al. 2020) pipeline for polarimetric data processing, and in total intensity using RDI through the DI-sNMF algorithm. The references were acquired via the “star-hopping” mode (Wahhaj et al. 2021), which means that the telescope is regularly pointed at a nearby reference star during data acquisition, efficiently alternating science and reference observations with minimal overhead. This observing strategy ensures maximal similarity between the respective PSFs.
@@ 5.1 Inspecting protoplanet claims in individual systems
Reggiani et al. (2014) and Biller et al. (2014) independently observed a PLF in $L^{\prime}$ band using independent ADI datasets from VLT/NACO processed with PCA. Better images of the disk obtained with SPHERE suggested that this candidate at $0\aas@@fstack{\prime\prime}11$ separation may rather trace filtered signal from the inner ring of the disk (e.g., Ligi et al. 2019). Later, another candidate was proposed using observations in polarized light from 2015, along with two ASDI observations in 2017 and 2019 using the IRDIS and multiple IFS observations between 2015 and 2019 (Gratton et al. 2019; Hammond et al. 2023). However, detecting the candidate with ASDI required suppressing most of the disk signal. The observations were conducted in $J$ and $YJH$ bands and the position of the candidate between 2015 and 2019 appeared to follow a Keplerian motion. The signal was interpreted as a gap-clearing Jovian-mass protoplanet surrounded by a circumplanetary disk or envelope. Studies by Garg et al. (2021) and Law et al. (2023) using ALMA observations revealed several distinct chemical signatures that demonstrate a compelling link to ongoing giant planet formation. Furthermore, compact CO J=2–1 and CO J=3–2 emission counterparts were observed that coincide with the location of the candidate planet. Yet, no PLF appears in polarized light nor in total intensity for both the DI-sNMF and IPCA-ARDI of the $K$-band images presented here, at its predicted location at the epoch of the Ks observations considered in this work (2021) assuming Keplerian motion. In Appendix F.2, we attempt to infer the detectability of the candidate by injecting a fake planet at the same separation, using the estimated contrast in Hammond et al. (2023). In this test, we processed the dataset with IPCA-ARDI as well as plain PCA-ADI. In addition, we provide the contrast curve obtained via PCA-ADI in Fig. 18. . Our $5\sigma$ detection limit of $1\times 10^{-4}$ at the expected location of the companion translates into $2M_{\text{Jup}}$ according to the ATMO2020 model, assuming no extinction. In comparison, Hammond et al. (2023) report a $YJH$ contrast of $\sim 1.5\times 10^{-5}$. Nevertheless, the detection of polarized intensity colocated with the protoplanet suggests the measured $YJH$ light is dominated by scattered light, making the mass conversion using atmospheric models most probably irrelevant. Assuming similar expected contrast in $Ks$ band as in $YJH$, our senitivity limits do not allow us to confirm or refute the candidate.
@@
Using CO line observations, Pinte et al. (2018) initially reported a local deviation from Keplerian velocity, suggesting the presence of a massive body at approximately 260 au from the star. Subsequently, Izquierdo et al. (2022) reobserved the first candidate proposed by Pinte et al. (2018), along with a second localized deviation from Keplerian rotation, potentially associated with the presence of a giant planet at 94 au. The first candidate proposed by Pinte et al. (2018) and Izquierdo et al. (2022) at 260 au does not fit within the field of view of the images presented in Figs. 11 and 9. At these separations, the motion of the candidates between the date of the first detection and the new dataset is negligible. We show another post-processed image for the same dataset, with an uncropped field-of-view in Appendix F.4 as well as contrast curve with PCA-ADI. No PLF is observed at the locations of the two proposed sources in these direct imaging observations with both IPCA-ARDI and PCA-ADI, despite the estimated sensitivity reaching down to $3\times 10^{-4}$ in contrast (i.e., mass limit of $1M_{J}$) for the candidate at 260 au, and $6\times 10^{-6}$ in contrast (i.e., mass limit of $4M_{J}$) at 94 au (see details in Appendix F.4).
@@ 6 Conclusion
We also explored the problem of parameters’ optimization for IPCA algorithms. While a good choice of parameters is important, we found that multiple sets of parameters lead to similar results. Nevertheless, we identified different scenarios and matched them with specific parametrizations to provide optimal results in each case. For the brightest disks with a stable speckle field, it is better to start the iterative process at rank one and have multiple iterations per rank before increasing the rank (e.g., CPD-68 1894 or HD 36112 data presented in Sect. 5). This ensures that most of the self-subtraction effects have been mitigated before loosening the model of the speckle field by increasing the rank. Conversely, when the disk is faint relative to the speckle, it is necessary to start at a high rank and limit the number of iterations per rank to avoid the bright artifacts stemming from non-captured speckle field components from being propagated in the subsequent iterations (e.g., V1094 Sco or GM Aur datasets). Regarding the convergence of the iterative process, we acknowledge the difficulty in building a reliable stopping criterion, which is a common problem in computer science (Nocedal et al. 2002). It is particularly challenging in our case of application, as determining where to stop the iterative process involves choosing a good compromise between noise subtraction (non-propagation) as well as the filtering and recovery of extended signals (convergence), and also because different parts of the disk converge at different rates. Nevertheless, observing the convergence curves remains an interesting indicator of the optimal iteration to be reached.
@@
While a few planet candidates were reported via direct imaging among the sampled young stars, we did not reobserve any of these candidates. This is at least partly due to limitations in the achieved sensitivity in the considered filter ($Ks$), which renders the detection of these companions challenging or unfeasible. While confirmed detections of protoplanets within disks remain rare, with only a select number being reported in the literature (e.g., Keppler et al. 2018; Haffert et al. 2019), studying young systems in their formative stages embedding protoplanetary disks housing nascent planets is crucial for understanding the intricate process of planetary system formation. The application of IPCA-ARDI holds promise in providing valuable insights by preserving both the structures of the disk and clear images of planets.
@@ Appendix A Local convergence of IPCA-ARDI
Figure 12: Evolution of the IPCA with ARDI estimate through the iterations. The dataset used is the same as in Fig. 3. The first grid of plots (top) displays the final estimation segmented into patches of size $15\times 15$ pixels. The second grid of plots (bottom) corresponds to the gradient computed within these patches according to Eq. 4. The red lines indicate the $g_{\rm tol}$ threshold at ±$\pm 10^{-2}$. If the threshold is reached, the curves are displayed in green, and a black dot indicates at which iteration the signal has converged. Otherwise, the curves are displayed in blue.
In this appendix, we study the evolution of the integrated flux into different patches of the image through the IPCA iterative process when using ARDI. Figure 12 shows that the brightest part of the disk converges faster than the rest of the image. We can also observe that the parts of the image where no signal is present do not seem to converge. It is important to note that the observed non-convergence on the outer parts of the image where there is no disk signal relates to the expression of the gradient. Indeed, from Eq. 4 it is clear that for an estimate $d_{i}$ that tends toward zero, the quantity $\frac{d_{i}-d_{i+1}}{d_{i}}$ becomes more sensitive to minor variations.
@@ Appendix C Comparison of results for different metrics
Figure 14: Same as Fig. 6 but for the four different metrics from top to bottom, respectively: Spearman’s rank, Pearson correlation, Euclidean distance, and SAD.
@@ Appendix D Results of disk estimations
Table 2: IPCA parameters used to reprocess each dataset from Ren et al. (2023) in Sect. 5. For each dataset, this table provides the starting rank and the number of iterations per rank. Regarding the number of iterations per rank, “incr” stands for incremental and indicates that the number of iterations increases by one at each rank (e.g., one iteration is performed at the starting rank $q$, two at rank $q+1$, and so on). Additionally, we also provide the number of iterations performed to obtain the final frame selected from visual assessment (“final iter.”), the rank that was reached at this iteration (“final rank”), and the mean and standard deviation of the PCC between science images and references.
@@ F.1 HD 36112 (MWC 758)
Figure 15: Results of IPCA (left) and 3-rank PCA with ADI only (right) for the VLT/SPHERE dataset of HD 36112 (Ren et al. 2023). Three simulated companions at a separation of $0\aas@@fstack{\prime\prime}6$ have been injected into the cube, with contrasts of $8\times 10^{-6}$, $1\times 10^{-5}$, and $2\times 10^{-5}$. White lines indicate their positions. Two squares are displayed to mark the locations of the two claimed protoplanets: a red square at $0\aas@@fstack{\prime\prime}1$, indicating the candidate proposed by Reggiani et al. (2018), and a yellow square at $0\aas@@fstack{\prime\prime}6$, referring to the candidate proposed by Wagner et al. (2019).
Figure 16: Contrast curve at $5\sigma$ confidence for the HD 36112 dataset, processed using rank-3 PCA with ADI only.